OpenStudy (anonymous):

What is the solution to the rational equation 4x/3 divided by 1/3=5?

OpenStudy (lgbasallote):

\[\LARGE \frac{\frac{4x}{3}}{\frac{1}{3}} = 5\] hint: you can cancel 3

OpenStudy (anonymous):

yah